OpenStudy (anonymous):

someone please help! ∫sin^3x / 1-cosx dx

terenzreignz (terenzreignz):

Well, this is a nice substitution problem... \[\Large \int \frac{\sin^3(x)}{1-\cos(x)}dx\]

terenzreignz (terenzreignz):

Why don't you let u = cos(x) and work from there? :)

OpenStudy (anonymous):

i think is better if u multiply with (1+cos(x))/(1+cos(x)) and see what u get :)

OpenStudy (anonymous):

du^3 / 1- u???

terenzreignz (terenzreignz):

whoa now... let's not get too far :P

terenzreignz (terenzreignz):

Here, let me rearrange it for you.. \[\Large=-\int \frac{[\sin^2(x)][-\sin(x)]}{1-\cos(x)}dx\]

OpenStudy (anonymous):

i tried multiplying both sides you get sin^3 x + sin^3x cosx all divided by sin 2 and then im stuck with a sinxcosx after simplifying it

terenzreignz (terenzreignz):

So, letting u = cos(x) du = -sin(x) dx

terenzreignz (terenzreignz):

Work from there?

OpenStudy (anonymous):

i thought the derivative of cosx is sinx ?

OpenStudy (anonymous):

if you multiply with( 1+cosx)/(1+cosx) you get (sin^3x*(1+cosx))/(1-cos^2x) then you have (sin^3(x)(1+cos(x)))/sin^2(x) , then its easy

terenzreignz (terenzreignz):

No... the INTEGRAL of cos(x) is sin(x) The derivative of cos(x) is -sin(x)

OpenStudy (anonymous):

oh ok but if i sub it into the modified equation i get -du^2 +du / 1 - u

OpenStudy (anonymous):

\[\int\limits\frac{ \sin ^{3} x }{1-\cos x }dx=\int\limits \frac{\sin ^{3}x }{ 1-\cos x }*\frac{ 1+\cos x }{ 1+\cos x }dx\] \[=\int\limits \frac{ \sin ^{3}x \left( 1+\cos x \right) }{ 1-\cos ^{2}x }dx\] \[=\int\limits \frac{ \sin ^{3}x \left( 1+\cos x \right) }{\sin ^{2}x}dx\] \[=\int\limits \left( \sin x+\sin x \cos x \right)dx=-\cos x+\frac{ \sin ^{2}x }{ 2 }+c\]

OpenStudy (anonymous):

@surjithayer thats so helpful but how did you get the integral of sinxcosx

OpenStudy (anonymous):

\[\int\limits f ^{n }\left( x \right)f'\left( x \right)dx=\frac{ f ^{n+1}\left( x \right) }{ n+1 }\]